Kelly's sister is 5 years younger than double Kelly's age. If the sum of their ages is sixteen how old is Kelly's sister?
Answer by stanbon(75887)

Kelly's sister is 5 years younger than double Kelly's age. If the sum of their ages is sixteen how old is Kelly's sister?
----
Equations:
s = 2k-5
s + k = 16
------
Substitute for "s" and solve for "k:
2k-5+k = 16
3k = 11
kelly = 11/3 3 yrs 8 months
-----------------------------
Solve for "s":
s = 2k-5
s = 22/3 - 15/3
--
sister's age = 7/3 = 2yrs 4 months
------------------------------
Cheers,
Stan H.